A person invests some money at 14%, with $8,000 more than this amount invested at 12%. The total annual income from the investments is $9,280.
How much is invested at 14%?

Answer by Boreal(15235)

X=INVESTED AT 14%
X+8000=INVESTED AT 12%
INCOME=
.14X+(.12)(x+8000)=.14X+.12X+960=9280
.26X=8320
X=$32000 answer
X+$8000=$40000
INTEREST ON FIRST IS $4480
(.12)(x+8000)=$4800
CHECKS